OpenStudy (anonymous):

please help? simplify sqrt-49 -------- (3+4i)-(2-5i)

OpenStudy (anonymous):

you need the sqaure root of -49?

OpenStudy (anonymous):

@Japs here, this is what my problem looks like

OpenStudy (anonymous):

i believe the answer is 7 i, i think i hope it helps and is right. :D

OpenStudy (anonymous):

\(\sqrt{-49}=\sqrt{49}\times \sqrt{-1}=7i\)

OpenStudy (anonymous):

\((3+4i)-(2-5i)=3+4i-2+5i=1+9i\)

OpenStudy (anonymous):

@satellite73 these arent helping?

OpenStudy (anonymous):

that is what you need to start the problem

OpenStudy (anonymous):

you need to start with \[\frac{7i}{1+9i}\] and then write in standard from by multiplying top and bottom by \(1-9i\)

OpenStudy (anonymous):

so the bottom part is 10 with the who equation as a negative?

OpenStudy (anonymous):

@satellite73?

OpenStudy (anonymous):

@Japs thanks!